OpenStudy (anonymous):

Solve the triangle. A = 50°, b = 13, c = 6

OpenStudy (danjs):



OpenStudy (anonymous):

Yes I get that but how do I find my missing pieces

OpenStudy (danjs):

You can use the law of sines or cosines

OpenStudy (anonymous):

I don't get how to do that I'm going virtual school and I need someone to walk me step by step

OpenStudy (danjs):

\[a^2 = 6^2 + 13^2 - 2(6)(13)*\cos(50)\]

OpenStudy (anonymous):

I don't have a calculator except my phone so when I do it, I don't get any of my answer choices

OpenStudy (danjs):

you get side a from that one, then you can use law of sines to get the last 2 angles \[\frac{ \sin(A) }{ a } = \frac{ \sin(B) }{ b } = \frac{ \sin(C) }{ c }\]

OpenStudy (anonymous):

Is there anyway you can calculate that because my phone doesn't get me the correct answer

OpenStudy (danjs):

a^2 = 36 + 169 - 156*(0.6428) a^2 = 104.7 a = about 10.2

OpenStudy (danjs):

\[\frac{ \sin(50) }{ 10.2 }=\frac{ \sin(B) }{ 13 }\]

OpenStudy (danjs):

B= about 77.5 degrees

OpenStudy (danjs):

THe last angle is 180 - the other 2